# Production of dileptonic bound states in the Higgs boson decay

### v1

@@ V Numerical results and Conclusions
Theoretical study of rare exclusive decay processes of the Higgs boson represents an important task in the general direction of studying the Higgs sector in the Standard Model. In this work, the calculation of the observed Higgs boson decay widths is carried out within the framework of the relativistic quasipotential model, which we previously used in studying the pair production of mesons and baryons. The creation of bound states of particles in various reactions is also important for testing the theory of bound states in quantum field theory.
@@
In this work, we investigated the role of relativistic effects caused by the law of transformation of wave functions of bound states of particles during the transition to the rest frame, relativistic corrections in the decay amplitude itself and corrections to the wave function of positronium, dimuonium and ditauonium at zero. The leading corrections are of order $\alpha$ and $\alpha^{2}\ln\alpha$. In addition, along with relativistic corrections, we also took into account the effects of particle coupling in our numerical results in Table 1, understanding by a purely non-relativistic calculation result one in which the mass of the bound state of leptons is equal to the sum of the lepton masses. The bound state effects occur because various mass coefficients enter into the final result in large degrees. In the case of dimuonium and ditauonium, these corrections are of order of a percent, which is mainly due to the effect of vacuum polarization of order $O(\alpha)$ in the bound state wave function. For positronium, the leading corrections are those of order $O(\alpha^{2})$, which leads to their significant reduction to 0.01 percent.
In calculating decay widths, we examine the contribution of various decay mechanisms, the relative contribution of which changes in the transition from positronium to dimuonium and ditauonium. Numerous mass parameters $r_{i}$ in the presented analytical formulas for decay widths, which determine the final result, change greatly during the transition between lepton bound states. But we can definitely say that the contribution from bosonic and quark loops (see Fig. 7,8) has the greatest values, along with the contribution of the ZZ-mechanism in the case of pair leptonium production. The widths of rare Higgs boson decays were calculated earlier in [8]. In general, we can say that our results and [8] are consistent, although there are some differences in individual decay widths.

### v2

@@ V Numerical results and Conclusions
Theoretical investigation of rare exclusive decay processes of the Higgs boson represents an important task in the general direction of studying the Higgs sector in the Standard Model. In this work, the calculation of the observed Higgs boson decay widths is carried out within the framework of the relativistic quasipotential model, which we previously used in studying the pair production of mesons and baryons. The creation of bound states of particles in various reactions is also important for testing the theory of bound states in quantum field theory. The difference between the production of bound states of quarks and leptons is that in the case of production of leptons, all calculations, including various corrections, can be carried out analytically, since the wave function of the bound state is known in analytical form.
@@
In this work, we investigated the role of relativistic effects caused by the law of transformation of wave functions of bound states of particles during the transition to the rest frame, relativistic corrections in the decay amplitude itself and corrections to the wave function of positronium, dimuonium and ditauonium at zero. The leading corrections are of order $\alpha$, $\alpha^{2}\ln\alpha$ and $\alpha^{2}$. In addition, along with relativistic corrections, we also took into account the effects of particle coupling in our numerical results in Table 1, understanding by a purely non-relativistic calculation result one in which the mass of the bound state of leptons is equal to the sum of the lepton masses. The bound state effects occur because various mass coefficients enter into the final result in large degrees. In the case of dimuonium and ditauonium, these corrections are of order of a percent, which is mainly due to the effect of vacuum polarization of order $O(\alpha)$ in the bound state wave function. For positronium, the leading corrections are those of order $O(\alpha^{2}\ln\alpha,\alpha^{2})$, which leads to their significant reduction to 0.01 percent.
In calculating decay widths, we examine the contribution of various decay mechanisms, the relative contribution of which changes in the transition from positronium to dimuonium and ditauonium. Numerous mass parameters $r_{i}$ in the presented analytical formulas for decay widths, which determine the final result, change greatly during the transition between lepton bound states. But we can definitely say that the contribution from W-boson and quark loops (see Fig. 7,8) has the greatest values, along with the contribution of the ZZ-mechanism in the case of leptonium production. The widths of rare Higgs boson decays were calculated earlier in [8]. In general, we can say that our results and [8] are consistent, although there are some differences in individual decay widths. These differences may be related to the different production mechanisms of lepton bound states investigated in our work.